• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:03:08 on October 9, 1998

    Entry number 2539

    keyword=Beam Time Accounting
    Hall A

    10/9/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: s incerti
    Shift Workers: L Todor,, J LeRose, r gilman

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0 0 0 0 0 0 0.5 1 1.5
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    1 1 1 1 1 1 0.5 0 6.5
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 0.6 0.4 1 1 1 7
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0.4 0.6 0 0 01
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 08:00 AM : day shift on standby 10:05 AM : target polarization ~ 28 % 10:35 AM : problems to start a run. Rebooted Coda and components 10:50 AM : started a cosmic run (#1640) 11:35 AM : target polarization ~ 29 % 11: 45 AM - 12:20 PM : access (cf Ron's note) 1:30 PM : Moeller quad tripped 3:00 PM : start BCM calibration

    Data Entry by: incerti